# Inferring Pluggable Types with Machine Learning

### v1

@@
Abstract
Pluggable type systems allow programmers to extend the type system of a programming language to enforce semantic properties defined by the programmer. Pluggable type systems are difficult to deploy in legacy codebases because they require programmers to write type annotations manually. This paper investigates how to use machine learning to infer type qualifiers automatically. We propose a novel representation, NaP-AST, that encodes minimal dataflow hints for the effective inference of type qualifiers. We evaluate several model architectures for inferring type qualifiers, including Graph Transformer Network, Graph Convolutional Network and Large Language Model. We further validated these models by applying them to 12 open-source programs from a prior evaluation of the NullAway pluggable typechecker, lowering warnings in all but one unannotated project. We discovered that GTN shows the best performance, with a recall of .89 and precision of 0.6. Furthermore, we conduct a study to estimate the number of Java classes needed for good performance of the trained model. For our feasibility study, performance improved around 16k classes, and deteriorated due to overfitting around 22k classes.
Index Terms:
component, formatting, style, styling, insert
I Introduction
@@
When programmers write code, they are often aware of semantic properties that a variable must adhere to but which are not checked by the type system of their programming language. Pluggable type systems are a method of adding custom type qualifiers to a language. A pluggable typechecker statically checks these user-defined qualifiers and warns about potential violations at compile time. For example, PUnits [1] adds qualifiers for scientific units, while Uber’s NullAway [2], Meta’s NullSafe [3], and the Checker Framework [4, 5] enforce programs do not crash due to null-pointer dereferences with a @Nullable qualifier. A key barrier to bringing legacy codebases into compliance with a typechecker is that developers must write these qualifiers into their code to see the full benefit of the tools. The state of the practice when deploying a new typechecker on a legacy codebase remains manual annotation: that is, in practice, type qualifiers are written by hand. The high cost of this manual process discourages developers from using these systems, despite their promise to improve code quality. This paper addresses the problem of bringing a legacy codebase into compliance with a typechecker by automatically inferring type qualifiers.
@@
Recently, researchers have proposed using machine learning to solve a similar problem: inferring type annotations for untyped Python or TypeScript programs [6, 7, 8, 9, 10, 11]. This problem shares a key feature with our problem of inferring type qualifiers for a legacy codebase: in both cases, the programs were written without feedback from the typechecker, so we expect the programs not to typecheck cleanly (with any set of annotations). Experimental results suggest that these techniques outperform traditional deductive type inference techniques, so in this paper we investigate whether a similar approach can work for the problem of inferring pluggable types.
@@
However, techniques for inferring types in Python or JavaScript programs are not immediately transferrable to our setting, because of three key differences. First, when inferring types for Python or JavaScript, the property of interest is known in advance: the type system is fixed (it enforces “tradtional” type safety: objects do not try to perform invalid operations). A key advantage of pluggable type systems is their customizability: each type system enforces its own rules. So, we require a way to customize our inference approach that does not require us to start over from scratch for each new type system. Second, for a given pluggable type system, the space of possible type qualifiers is fixed: for example, a typechecker for preventing null dereferences might only contain the qualifiers @NonNull and @Nullable. By constrast, when inferring types for e.g., Python, the space of possible types is infinite and has a long tail, because of the presence of user-defined types. That is, in the standard setting the type system is fixed and the number of types is infinite; but in the pluggable types setting there are infinitely-many possible type systems, but each has a finite, known universe of types. Third, machine learning is data hungry: approaches for popular languages like Python can assume the existence of a large corpus of human-written types in open-source software. By contrast, the customizability of pluggable types means that most type systems have few users: in fact, pluggable type systems are sometimes designed only for a single program! So, to apply machine learning techniques to the pluggable type inference setting, we need to understand how much data is actually required to make them effective. These differences inspired us to investigate three high-level research questions:
  • RQ1: What kind of ML model should we use, and how should we post-process its output?
  • RQ2: How should we encode the program in a type-system-specific way to train the model?
@@
  • RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?
@@
The state-of-the-art models [12, 6, 7] for inferring types in Python and JavaScript leverage the type dependence graph, a graph that encodes the type system’s semantics and which is therefore unsuitable for representing pluggable type systems because each type system has its own rules. Prior works used a variety of model architectures, including graph models and text-based models. For example, Typilus [12] learns a type space, STIR [6] uses a BiLSTM for incomplete code, and Type5 [7] uses a pre-trained seq2seq model. The differences between the pluggable types setting and the Python/JavaScript setting, however, inspired us to investigate RQ1 by evaluating a set of relatively-simple models. In particular, we eschew time-series models like BiLSTM (which are effective for sequential data, like recognizing the content of a video, but very expensive during training and inference) because we hypothesize that they may not be needed to predict pluggable type qualifiers. We considered two graph models and one text-based model. The two graph models are a baseline [13] graph convolutional network (GCN) model [14] and a more-capable graph transformer network (GTN) model [15], which extends the GCN’s capabilities by adding the ability to reason about multiple edge types (i.e., hypergraphs). The text-based model is a large language model (LLM) [16]; LLMs have performed well on coding tasks recently, but they operate in reduced context because of their token limit.
To investigate RQ2, we considered two techniques: 1) prompt the LLM about which program locations should be annotated with a type qualifier, and 2) develop a novel training methodology for building a type-system-specific graph encoding. The first technique was relatively ineffective, so we pursued the second. Our graph encoding methodology uses ablation studies to retain information that is useful for predicting the type qualifiers used by a particular type system. This process results in a novel program encoding that we call a Name-augmented, Pruned Abstract Syntax Tree (NaP-AST), which we describe in section III. A NaP-AST encodes the parts of the program’s structure that are relevant to the type system of interest, along with name information that ties related program elements together, into a graph that is a suitable input to a graph-based model.
@@
To evaluate these models and encodings, we conducted a feasibility study. Of the commonly used type qualifiers, those for nullability in Java are the most common. We therefore conducted our feasibility study on nullness: if there is not enough data for Java nullness type qualifiers to train an effective model, there is not enough data to train an effective model for any pluggable type system. First, we collected a dataset from Github of Java classes, each of which uses one of 31 different @Nullable annotations. After cleaning, this dataset contains 32,370 classes; more details about the dataset and its collection are presented in section IV. We then used our methodology for building a NaP-AST encoding to derive a NaP-AST encoding for nullness. Then, we trained the three models using this dataset: the two graph models (a GCN and a GTN) using the derived NaP-AST encoding, and the text-based LLM using syntax-based code chunking. We then compared their performance on 12 open-source projects that had previously been annotated for the popular NullAway typechecker [2]. The GTN outperformed the other models we evaluated: the GCN requires additional postprocessing to reach acceptable performance, while the LLM is too limited in token capacity to be effective on large projects. Our best model (the GTN) achieves 89% recall of the human-written type qualifiers and 60% precision; its inferred annotations eliminate 60% of the NullAway warnings that a human would need to triage (a proxy for human effort). Our success for nullability suggests that our approach may be effective with other type qualifiers, but unfortunately we lack enough data for any other type qualifier to test this hypothesis.
@@
However, to assist future researchers or practitioners who would like to train a similar model for another type qualifier (and answer RQ3), we explore how much training data one needs for adequate performance. We rebuilt our best model with random samples of the dataset and re-ran the evaluation to show how much data it needs before it acheives good performance. We found that around 16k human-annotated classes were needed: a large but not impossible-to-reach number. We counted 217,922 annotations in these classes.
The primary contributions of our paper are:
@@
  • A novel methodology to construct a per-type-system graph encoding of programs (NaP-ASTs) to train deep-learning graph models like GCNs and GTNs (section III);
@@
  • A dataset of 32,370 Java classes annotated with @Nullable that enables our feasibility study (section IV);
@@
  • A comparison of three model architectures (GCN, GTN, and LLM) that shows that the GTN is best for the problem of inferring pluggable types in our feasibility study (section VI-A), with open-source models and data;
@@
  • A study of the amount of data required by our best GTN model on the task of inferring type qualifiers for adequate performance (section VI-B).
II Background: Model Architectures
@@
This section provides background on the three considered models: graph convolutional networks (GCNs), graph transformer networks (GTNs), and large language models (LLMs). We focus on the model architectures and their salient features with respect to type qualifier inference, and also introduce the specific models that we use in our experiments.
II-A Graph Models
II-A1 GCN
@@
Graph Convolutional Networks (GCNs) are a class of neural networks designed for graph data. Type qualifier prediction can be modeled as a graph data task because ASTs can be represented as a graph. GCNs use graph convolution, a propagation rule that updates the feature representation of each node by aggregating features from its neighbors (and the node itself). For AST prediction, GCNs can be particularly effective due to explicitly modeling local context: ASTs are inherently hierarchical and structured, where the context around a node (e.g., its parent or children nodes) significantly influences it. GCNs leverage this local structure efficiently. They are also efficient in learning spatial features, which have been helpful for ASTs [17].
@@
II-A2 GTN
GCNs assume that the provided graphs are homogeneous, i.e., they consist of one type of nodes and edges. However, in uses cases like source code annotation distinct edge types can alter the importance of each node. Taking advantage of heterogenous graphs allows us to layer additional hints over bare ASTs. This paper uses it to add a layer connecting names across an AST, which provides dataflow hints.
@@ II-A2 GTN
To do so, we use the FastGTN model [18], which uses the meta-path, a multi-hop sequence of connected nodes and edges in the graph with heterogeneous edge types. Instead of ignoring the edge type information, breaking down a heterogeneous graph into homogeneous graph by its meta-paths allows the model to preserve this information. A Graph Transformer Network (GTN) [18] learns the most informative meta-paths automatically for each task in an end-to-end manner by selecting adjacency matrices through a 1x1 convolution with weights from a softmax function. The GTN can learn arbitrary meta-paths with respect to edge types and path length. GTNs consider multiple types of meta-paths simultaneously by generating multiple graph structures: they set the output channels of a 1x1 filter to the number of channels. After stacking $k$ GT layers, multi-layer GNNs are applied to each channel of the output tensor. The final node representations can be used for downstream tasks, by applying dense layers followed by a softmax layer to the node representations.
@@
A weakness of GTNs is that they are not scalable because of the high computational cost and memory requirements when calculating and storing new adjacency matrices at each layer. To address this, FastGTNs implicitly transform the graph structures without storing the new adjacency matrices of meta-paths, avoiding the need for explicit multiplications of large adjacency matrices. This reduces the computational cost and memory usage while keeping FastGTNs mathematically identical to GTNs. Even FastGTNs have serious scalability limitations, which motivates our pruning strategies discussed in section III; note that the primary scalability challenge for GTNs is fitting the $n$-by-$n$ adjacency matrix they require into memory (where $n$ is the size of the input)—even with “large” inputs, both training and prediction with FastGTNs is relatively quick (see further discussion in section VI-A3).
II-A3 GCN vs GTN
@@
The GTN is a generalization of a GCN that supports distinct edge types. This allows the GTN to distinguish the AST and the name layer. While GCNs do not automatically support this heterogeneous operation, they are a good baseline for studies exploring advanced graph learning techniques [13]. because they can integrate local neighborhood information through the graph convolution operation.
II-B Text-based Model: the Large Language Model (LLM)
Magicoder [16] is a fully open-source, state-of-the-art LLM trained on a dataset created from 75,000 synthetic instruction data generated using the OSS-INSTRUCT approach. OSS-INSTRUCT involves generating coding problems from real world snippets, ensuring diversity by incorporating a wide variety of challenges, and enhancing the controllability of the output. Since Magicoder has outperformed GPT 3.5 using only open source training samples, it is a suitable baseline for comparison with other techniques.
The LLM approach leverages a vast knowledge base, ensuring a diverse understanding of coding patterns and practices, significantly beneficial for inferring type qualifiers by understanding the broader context and nuances in code. LLMs have recently been used to great effect in a number of software engineering domains [19]. The model’s limitation is that if certain edge cases or less common programming constructs are underrepresented, the model’s inference capability might be constrained for those scenarios. Crucially, LLMs may struggle with deep semantic understanding of code, they are only predicting the next token in a sequence. In practice, there are other limits for using LLMs. For example, Magicoder limits its output to 2048 tokens.
III Encodings

### v2

# Nullability Type Inference via Machine Learning

@@
Abstract.
Pluggable type systems improve code safety by adding type qualifiers, but deploying them in legacy Java projects is difficult because annotations must be written by hand. This paper investigates using machine learning to infer type qualifiers. We evaluate graph models—Graph Convolutional Networks (GCNs) and Graph Transformer Networks (GTNs)—and large language models (LLMs). For graphs, we test Type Dependency Graph (TDG) encoding from prior work and introduce a novel encoding, Name-augmented, Pruned Abstract Syntax Tree (NaP-AST), embedding minimal dataflow hints. Models are validated on 9 open-source projects human annotated for the NullAway nullness checker. The GTN with NaP-AST achieves the best downstream impact, cutting checker warnings by 69%, with 69% recall of human-written annotations and 39% precision. The strongest LLM shows the opposite trade-off: 91% precision but only 21% recall, still reducing warnings by 61%. Regarding generalization, we conducted an experiment to estimate the number of classes required in the training set to achieve good performance: performance improves up to roughly 16K annotated classes and then degrades from overfitting around 22K.
pluggable types, machine learning, static analysis
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Conference Name; Month, Year; Locationbooktitle: Proceedings of the Conference Name (Conference Name ’24), Month, Year, Locationisbn: 978-1-4503-XXXX-X/24/08ccs: Software and its engineering
1. Introduction
@@
When programmers write code, they are often aware of semantic properties that a variable must adhere to but which are not checked by the type system of their programming language—for example, that a variable should only hold non-null values. Pluggable type systems (Foster et al., 1999; Bracha, 2004) add custom type qualifiers to a language. A pluggable typechecker statically checks these user-defined qualifiers and warns about potential violations at compile time. For example, PUnits (Xiang et al., 2020) adds qualifiers for scientific units, while Uber’s NullAway (Banerjee et al., 2019b), Meta’s NullSafe (Pianykh et al., 2022), and the Checker Framework’s Nullness Checker (Papi et al., 2008; Dietl et al., 2011) enforce programs do not crash due to null-pointer dereferences with an @Nullable qualifier. A key barrier to bringing legacy codebases into compliance with a pluggable typechecker is that developers must write these qualifiers into their code to see the full benefit. The state of the practice when deploying a new typechecker on a legacy codebase remains manual annotation: that is, type qualifiers are written by hand. The high cost of this manual process discourages developers from using these systems, despite their promise to improve code quality. This paper addresses the problem of bringing a legacy codebase into compliance with a null-safety typechecker by automatically inferring type qualifiers. While our primary focus is on nullability for data availability reasons, we propose a framework that is theoretically adaptable to other finite pluggable type systems, if sufficient training data were available.
@@
Recently, researchers have proposed using machine learning to solve a related, but different, problem: inferring type annotations for Python or JavaScript programs (Peng et al., 2023b; Wei et al., 2023; Peng et al., 2023a, 2022a; Mir et al., 2022; Cui et al., 2021). This problem shares a key feature with inferring type qualifiers for a legacy codebase: in both cases, the programs were written without feedback from the typechecker, so we expect them not to typecheck cleanly (with any set of annotations). Experimental results suggest that these techniques outperform traditional deductive type inference techniques. Moreover, large language models (LLMs) have recently shown promise on a variety of coding tasks (Roychoudhury et al., 2025; Miserendino et al., 2025; Mündler et al., 2024; Bouzenia et al., 2025; Yang et al., 2024; Ozkaya, 2023; Wu et al., 2023; Xiong et al., 2024; Ross et al., 2023). This paper’s goal is to investigate if similar approaches can work for inferring pluggable types.
@@
However, techniques for inferring types in Python or JavaScript programs are not immediately transferable to our setting, because of three key differences. First, when inferring types for Python or JavaScript, the type system is fixed (it enforces “traditional” type safety: objects do not try to perform invalid operations). A key advantage of pluggable type systems is their customizability: each type system enforces its own rules. So, we desire a way to customize our inference approach that does not require us to start over from scratch for each new type system. Second, when inferring types for e.g., Python, the space of possible types is infinite and has a long tail, because of the presence of user-defined types. Prior works focus on solving this so-called rare type problem; for example, by encoding programs using a Type Dependency Graph (TDG(Peng et al., 2022b) data structure which is specialized to help the model find rare types. However, while the tail of types is long, the set of locations that need a type annotation is not: in prior work, the locations where a type should be written are fixed in advance. For most pluggable type systems, these problems are reversed: the hardest problem is not choosing which type to write, but at which places to write a type. For example, for a nullability type system, an annotator usually only needs to write either @Nullable or @NonNull (and the pluggable type system interprets the lack of a qualifier as the other), so the problem that an inference tool needs to solve is where to write a qualifier rather than which type to write. This difference makes the encodings from prior works (e.g., the TDG) ill-suited to our setting. Third, machine learning is data hungry: approaches for popular languages like Python can assume the existence of a large corpus of human-written types in open-source software. The customizability of pluggable types is a double-edged sword: many pluggable type systems are highly-specialized and therefore have relatively few users. So, to apply machine learning techniques to the pluggable type inference setting, we need to understand how much data is actually required to make them effective. These differences inspired us to investigate two high-level research questions:
  1. RQ1
    What kind of machine-learning model architecture is suitable for inferring pluggable types, and how does that model differ from those in prior type inference works for other languages?
  2. RQ2
@@
    How much data is needed to make that model effective at inferring qualifiers in practice?
@@
State-of-the-art models (Peng et al., 2023b; Wei et al., 2023) for inferring types in Python/JavaScript use diverse model architectures, including graph and text-based models: Typilus (Allamanis et al., 2020) learns a type space, STIR (Peng et al., 2023b) uses a BiLSTM for incomplete code, and TypeT5 (Wei et al., 2023) uses a pre-trained seq2seq model. The differences with the pluggable types setting, however, inspired us to investigate RQ1 by evaluating a set of models used in previous work. We eschew time-series models like BiLSTM (which are effective for sequential data, like recognizing the content of a video, but very expensive during training and inference) because predicting where to write a type qualifier is more similar to a classification task. We considered four models: two graph models and two text-based models. For the graph models, we considered two encodings: the TDG encoding (Peng et al., 2022b) from prior work and a novel encoding that we developed specifically for pluggable type inference. Our novel graph encoding methodology uses ablation studies to retain information useful for predicting the type qualifiers of a particular type system, which results in a novel program encoding that we call a Name-augmented, Pruned Abstract Syntax Tree (NaP-AST), which we describe further in section 3. A NaP-AST encodes the parts of the program’s structure that are relevant to the type system of interest, along with name information that ties related program elements together, into a graph that is a suitable input to a graph-based model. The first graph model is a baseline (Wu et al., 2019) graph convolutional network (GCN) model (Zhang et al., 2019a). The second is a more-capable graph transformer network (GTN) model (Yun et al., 2019), which extends the GCN to reason about multiple edge types (i.e., hypergraphs). Since the TDG encoding is not a heterogenous graph, we only use it with the GCN model; the NaP-AST encoding is a heterogenous graph, so it can benefit from the GTN model. Both text-based models are large language models (LLMs), which we include because LLMs have performed well on coding tasks recently. The first is CodeQwen1.5-7B-Chat  [^1] https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat, an open-source model specialized for coding tasks. The second text-based model is GPT-4 (Achiam et al., 2023), a state-of-the-art [^2] GPT-4 was OpenAI’s flagship model when these experiments were conducted, though at the time of submission OpenAI recently (August 7, 2025) released its new GPT-5 model. commercial LLM.
@@
To evaluate these models and encodings, we conducted a feasibility study: we trained models to predict where to place type qualifiers for one popular pluggable type system. We would have preferred to repeat this experiment for more than one type system, to show the generality of our approach, but a lack of publicly-available data prevents us from doing so; we show this empiricially in section 7.3.1. Of the commonly used type qualifiers, those for nullability in Java are the most common. We therefore conducted our feasibility study on nullness: if there is not enough data for Java nullness type qualifiers to train an effective model, there is not enough data to train an effective model for any pluggable type system. First, we collected a dataset from GitHub of Java classes, each of which uses one of 31 different @Nullable annotations. After cleaning, this dataset contains 32,370 classes. We then used our methodology for building a NaP-AST encoding to derive a NaP-AST encoding for nullness. Then, we trained the four models using this dataset: the two graph models (a GCN and a GTN) using the derived NaP-AST encoding, the GCN using the baseline TDG encoding (which cannot benefit from the GTN, because it is a homogenous graph), and the text-based LLMs (CodeQwen1.5-7B-Chat and GPT-4) using syntax-based code chunking. We then compared their performance on 9 open-source projects that had previously been annotated for the popular NullAway typechecker (Banerjee et al., 2019b). The GTN with the NaP-AST encoding outperformed the other models we evaluated: both GCN models have lower recall than NullGTN; TDG GCN has higher precision than NullGTN but fails to reduce warnings, while NullGCN trails on both precision and recall. LLMs can hallucinate edits at scale; CodeQwen attains very high precision (0.91) but low recall (0.21), and GPT-4 has high precision (0.76) with moderate recall (0.25). Our best model (which we call NullGTN) achieves 69% recall of the human-written type qualifiers and 39% precision; its inferred annotations eliminate 69% of the NullAway warnings that a human would need to triage (a proxy for human effort).
@@
Our success for nullability suggests that our approach may be effective with other type qualifiers, but unfortunately we lack enough data for any other type qualifier to test this hypothesis. To assist future researchers or practitioners who want to train a similar model for another type system (and to answer our RQ2), we explore how much training data one needs for adequate performance. We rebuilt our best model with random samples of the dataset and re-ran the evaluation to show how much data it needs before it achieves good performance. Around 16k human-annotated classes were needed: a large but not impossible-to-reach number. Our primary contributions are:
@@
  • A novel methodology to construct a graph encoding of programs for null safety (NaP-ASTs) to train deep-learning graph models like GCNs and GTNs that can theoretically be extended to other finite type systems (section 3);
@@
  • A dataset of 32,370 Java classes annotated with @Nullable that enables our feasibility study (section 4);
@@
  • A comparison of three model architectures (GCN, GTN, and LLM) and two graph encodings (NaP-AST and TDG) that shows that the GTN with the NaP-AST encoding is best at inferring pluggable types in our feasibility study (section 7.1), with open-source models and data;
@@
  • A study of the amount of data required by our best GTN model on the task of inferring type qualifiers for adequate performance (section 7.3).
2. Background: Model Architectures
@@
This section describes the three considered model architectures: graph convolutional networks (GCNs), graph transformer networks (GTNs), and large language models (LLMs). We focus on their salient features with respect to type qualifier inference.
2.1. Graph Models
2.1.1. GCN
@@
Graph Convolutional Networks (GCNs) are a class of neural networks for graph data. Type qualifier prediction can be modeled as a graph data task since ASTs can be represented as a graph. GCNs use graph convolution, a propagation rule that updates the feature representation of each node by aggregating features from its neighbors and the node itself. For AST prediction, GCNs can be particularly effective, since they explicitly model local context: ASTs are inherently hierarchical and structured, where the context around a node (e.g., its parents or children) significantly influences it. GCNs leverage this local structure efficiently. They are also efficient in learning spatial features, which have been helpful for ASTs (Zhang et al., 2019b).
@@
2.1.2. GTN
GCNs assume their input graphs are homogeneous, i.e., they consist of one type of nodes and edges. Using a model that supports heterogenous graphs lets us layer hints over the AST, which we use to add a layer connecting names across an AST to provide dataflow hints.
@@ 2.1.2. GTN
We use the FastGTN model (Yun et al., 2022), which uses the meta-path, a multi-hop sequence of connected nodes and edges in the graph with heterogeneous edge types. Instead of ignoring the edge type information, breaking down a heterogeneous graph into homogeneous graph by its meta-paths allows the model to preserve this information. A Graph Transformer Network (GTN) (Yun et al., 2022) learns the most informative meta-paths automatically for each task in an end-to-end manner by selecting adjacency matrices through a 1x1 convolution with weights from a softmax function. The GTN can learn arbitrary meta-paths with respect to edge types and path length. GTNs consider multiple types of meta-paths simultaneously by generating multiple graph structures: they set the output channels of a 1x1 filter to the number of channels. After stacking $k$ GT layers, multi-layer GNNs are applied to each channel of the output tensor. The final node representations can be used for downstream tasks, by applying dense layers followed by a softmax layer to the node representations.
MethodDeclSimpleName”f”SimpleNameMethodCallExprAST ChildName ParentName ChildAST Parent
Figure 1. In heterogeneous graphs (i.e. having multiple edge types), a meta-path is a sequence of connected nodes and edges, where each edge has a distinct type.
@@
A weakness of GTNs is that they are not scalable because of the high computational cost and memory requirements when calculating and storing new adjacency matrices at each layer. To address this, FastGTNs implicitly transform the graph structures without storing the new adjacency matrices of meta-paths, avoiding the need for explicit multiplications of large adjacency matrices. This reduces the computational cost and memory usage while keeping FastGTNs mathematically identical to GTNs. Even FastGTNs have serious scalability limitations, which motivates our pruning strategies discussed in section 3; note that the primary scalability challenge for GTNs is fitting the $n$-by-$n$ adjacency matrix they require into memory (where $n$ is the size of the input)—even with “large” inputs, both training and prediction with FastGTNs is relatively quick (see further discussion in section 7.2).
2.1.3. GCN vs GTN
@@
The GTN is a generalization of a GCN that supports distinct edge types. This allows the GTN to distinguish the AST and the name layer. While GCNs do not automatically support this heterogeneous operation, they are a good baseline for studies exploring advanced graph learning techniques (Wu et al., 2019) because they can integrate local neighborhood information through the graph convolution operation.
2.1.4. Models from Prior Work
Although we did test the TDG encoding (section 3.1) from prior work, we did not evaluate the Python-specific models with which it was originally deployed, because they are not transferable to our setting. In particular, the Typilus paper (Allamanis et al., 2020) used a custom “TypeSpace”, metric-based meta-learning, and a Graph Neural Network (GNN) model optimized for predicting rare types. None of this machinery is sensible or necessary for predicting pluggable type qualifiers, which are typically drawn from a fixed set—for example, in our feasibility study, from the set containing @Nullable and @NonNull.
2.2. Text-based Models
2.2.1. CodeQwen1.5-7B-Chat
CodeQwen 1.5-7B-Chat [^3] https://qwenlm.github.io/blog/codeqwen1.5/ is a fully open-source transformer (7 B params, 64 K context) trained on 3 T code tokens spanning 92 programming languages, and attains 83.5% on HumanEval-plus and 78.7% on MBPP-plus, outscoring other 7B code models  [^4] https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat. The model is fine-tuned with a synthetic instruction corpus that couples long-context code snippets with conversational tasks (Bai et al., 2023). Because CodeQwen 1.5-7B-Chat outperforms StarCoder 2, CodeLlama 7 B and DeepSeek-Coder on major code benchmarks, it is an appropriate open-source baseline. The LLM approach leverages a vast knowledge base, ensuring a diverse understanding of coding patterns and practices, significantly beneficial for inferring type qualifiers by understanding the broader context and nuances in code. LLMs have recently been used to great effect in a number of software engineering domains (Roychoudhury et al., 2025; Miserendino et al., 2025; Mündler et al., 2024; Bouzenia et al., 2025; Yang et al., 2024; Ozkaya, 2023; Wu et al., 2023; Xiong et al., 2024; Ross et al., 2023). Their limitation is that they may struggle with deep semantic understanding of code, as they are only predicting the next token in a sequence. Specifically, LLMs hallucinate about code structure. These hallucinations add up when scaled to hundreds of thousands of lines of code. While CodeQwen benefits from the 64 K-token window, it can still hallucinate nonsensical changes, like adding or removing import statements in large files. All current LLM systems hallucinate since their function is to predict the next token. To get around this limitation, we only asked the LLM “yes or no” questions, restricting its ability to change code (see section 3.3 for more details). In practice, there are other limits on LLMs, like token limits and cost.
2.2.2. GPT-4o-mini
GPT-4o-mini is an LLM that demonstrates exceptional performance in coding and natural language processing tasks, outperforming its predecessors and many contemporary models in various domains. For example, GPT-4o-mini achieves superior accuracy in programming competitions and exhibits strong cross-language translation capabilities, showcasing its adaptability and generalization skills (Hou and Ji, 2024). Its ability to refactor and annotate code with significant quality improvements further highlights its utility in software engineering (Poldrack et al., 2023).
Although excelling in generalization, GPT-4o-mini still faces notable limitations. Its reliance on token limits restricts its ability to process large datasets in a comprehensive way, and exhibits a tendency to hallucinate or produce errors in domain-specific contexts, such as radiology or highly technical programming tasks (Liu et al., 2023; Pordanesh and Tan, 2024). Like with CodeQwen, we restricted its ability to change code, asking only yes or no questions.
3. Encodings
serialize()bytes (byte[])has_parameter
Figure 2. An example TDG snippet. A function $serialize()$ that has a parameter called bytes of type $byte[]$. $serialize()$ is an expression node. bytesbytes is a symbol node. $Has_{p}arameter$ is a merge node.
This section details the three encodings that we considered. The first two encodings are for the graph model architectures (GCNs and GTNs); the last encoding is for the text-based LLM models:
  • the type dependency graph (TDG) encoding from prior work (section 3.1),